To start an avalanche on a mountain slope, an artillery shell is fired with an initial velocity of 310 m/s at 57.0° above the horizontal. It explodes on the mountainside 37.0 s after firing. What are the x and y coordinates of the shell where it explodes, relative to its firing point?

x=? (m)

y=? (m)
Answer #1

u= 310m/s, g- 570 ㄐ ㄧ-.. , (Y hooil me tho = 210 cos(57) X 37
Answer #2

SOLUTION :



Shell is fired at an angle of 57º and speed of 310 m/s. 

upward. It travels for 37 seconds and explodes.


So, 


horizontal speed  of shell , Vh = 310 cos (57) =  168.8381 m/s


vertical speed  of shell  , Vv = 310 sin (57) = 259.9879  m/s


Coordinates of explosion point relative to firing point(0, 0) :


X = Horizontal distance moved by shell = Horizontal speed * time 

= 168.8381 * 37 = 6247.00 meters 


Y = vertical distance travelled by shell 

= u t - 1/2 g t^2   ;  (u = Vv)

= (259.9879)(37) - 1/2 * 9.8 * (37)^2 = 2911.45  meters.


Hence, X, Y coordinates of explosion point are : 6247.00 m and 2911.45 m respectively.

(ANSWER).



answered by: Tulsiram Garg